if xylitol has a chemical formula of c5h12o5, what is the percentage of oxygen in the compound?

1
Expert's answer
2022-02-01T04:58:04-0500

Solution:

Xylitol = C5H12O5


Atomic mass of elements:

Ar(C) = 12.0107 amu

Ar(H) = 1.00784 amu

Ar(O) = 15.999 amu


Calculate the molar mass of xylitol (C5H12O5):

M(C5H12O5) = 5×Ar(C) + 12×Ar(H) + 5×Ar(O)

M(C5H12O5) = 5×12.0107 + 12×1.00784 + 5×15.999 = 152.1426 (g/mol)


Calculate the percentage of oxygen in the compound:

w(O) = [5×Ar(O) / M(C5H12O5)] × 100%

w(O) = [5×15.999 / 152.1426] × 100% = 52.58%

w(O) = 52.58%


Answer: The percentage of oxygen (O) in xylitol (C5H12O5) is 52.58%